OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 3x

OpenStudy (anonymous):

cos(3x) = cos(x+x+x) = cos(x+2x) = cos(x)*cos(2x) - sin(x)*sin(2x) = cos(x)*(cos(x)*cos(x) - sin(x)*sin(x)) - sin(x)*(sin(x)*cos(x)+cos(x)*sin(x)) =cos^3(x) - sin^2(x)*cos(x) - sin^2(x)*cos(x) - sin^2(x)*cos(x) = cos^3(x) - 3*sin^2(x)*cos(x)

OpenStudy (anonymous):

um the answer choices are cos x - 4 cos x sin2x -sin3x + 2 sin x cos x -sin2x + 2 sin x cos x 2 sin2x cos x - 2 sin x cos x

OpenStudy (anonymous):

i assumed right

OpenStudy (anonymous):

cos x - 4 cos x sin^2x -sin^3x + 2 sin x cos x -sin^2x + 2 sin x cos x 2 sin^2x cos x - 2 sin x cos x

OpenStudy (anonymous):

@DUDE..IM..A..DUCK Then how do I do it? btw don't just give the answer

OpenStudy (anonymous):

mind if i help?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[\cos(3x)=\cos(x+2x)\]\[\cos(x+2x) = (cosx)(\cos2x) - (sinx)(\sin2x)\]\[=(cosx)(\cos^{2}x-\sin^{2}x) - sinx(2sinxcosx)\]\[=cosx(\cos^{2}x-\sin^{2}x-2\sin^{2}x)\]

OpenStudy (anonymous):

do you follow me up till here?

OpenStudy (anonymous):

so far yes

OpenStudy (anonymous):

since\[\cos^{2}x = 1-\sin^{2}x\]\[cosx(1-\sin^{2}x-\sin^{2}x-2\sin^{2}x)\]\[cosx(1-4\sin^{2}x)\]\[cosx-4cosxsin^{2}x\]

OpenStudy (anonymous):

and you are done...

OpenStudy (anonymous):

ask me any questions if you dont get something

OpenStudy (anonymous):

wow thanks but how did you get this or do this part? =(cosx)(cos^2x−sin^2x)−sinx(2sinxcosx) =cosx(cos^2x−sin^2x−2sin^2x)

OpenStudy (anonymous):

I am just taking out cosx as the factor

OpenStudy (anonymous):

i will do that part in more detail... please wait

OpenStudy (anonymous):

ok no problem

OpenStudy (anonymous):

\[cosx(\cos^{2}x-\sin^{2}x)-sinx(2sinxcosx)\]\[cosx(\cos^{2}x-\sin^{2}x) - 2\sin^{2}xcosx\]\[cosx(\cos^{2}x-\sin^{2}x-2\sin^{2}x)\]

OpenStudy (anonymous):

do you see now?

OpenStudy (anonymous):

okay but real quick what happend to the xcosx?

OpenStudy (anonymous):

there is no xcosx... I think you are mistaking (sinx)(cosx), normally writtenas sinxcosx with sin xcosx...

OpenStudy (anonymous):

oh okay got it! haha thanks

OpenStudy (anonymous):

haha... have fun and good luck with similar questions...

OpenStudy (anonymous):

thanks